# Bid Farewell to Seesaw: Towards Accurate Long-tail Session-based Recommendation via Dual Constraints of Hybrid Intents

### v1

@@ Introduction
Figure 1: Comparison between the our proposed HID and previous work. (a) illustrates the design of HID, where t. and n. denotes target and noise items for session $S^{u}$, respectively; (c) and (d) demonstrate the frameworks of previous long-tail approaches. (b) evaluates the accuracy (i.e., HR@20) and long-tail performance (i.e., tCov@20) of the base SBR model GRU4Rec (Hidasi16gru4rec) and GRU4Rec + long-tail approaches on Tmall dataset.
@@
Session-based recommendation (SBR) addresses information overload by predicting the next item from short-term interactions, particularly in privacy-sensitive scenarios lacking long-term user profiles (Li2025Survey; Latifi21Session). While deep learning methods in SBR (e.g., deep sequential models (Hidasi16gru4rec; Li17Narm; liu18Stamp; yuan21Dsan; Hou22Core) and deep graphic models (Wu19Srgnn; Qiu19Rethinking; Wang20Gcegnn; Xia21Dhcn; Xia21Cotrec; Pan20Star)) can effectively model item correlations, their model-centric focus overlooks inherent data biases. A key challenge is the long-tail distribution in recommendation data (Sundaresan2011Recommender; Yang2023LOAM; Liu2020TailNet), where a small number of high-exposure items (i.e., head items) dominate the model’s attention, while a significantly larger number of low-exposure items (i.e., tail items) are often disregarded. This unfair phenomenon leads to the overlooking of potentially essential but low-exposure tail items, limiting the diversity of recommendations (Turgut2023Diversity; Yin2024Diversity; Lee2024PostTraining). Besides, the long-tail distribution causes the model to be more inclined to recommend head items, resulting in a vicious cycle.
@@
Previous advancements in long-tail SBR focus on developing plugins that seamlessly integrate with existing SBR models, emphasizing the significance of tail items to mitigate the long-tail issue (Liu2020TailNet; Chen2023CSBR; Yang2023LOAM; Peng2024LAPSR). Broadly, long-tail SBR approaches fall into two categories: (i) Augment-based approaches, which employ augmentation strategies to refine the tail item embeddings or session embeddings (Yang2023LOAM; Kim2023MELT; Huang2024CSLP; Liu24LLMESR), and (ii) Rerank-based approaches, which predict head/tail item distributions based on interaction sessions and directly modify the final ranking results (Liu2020TailNet; Chen2023CSBR; Peng2024LAPSR). Both approaches consistently emphasize the significance of tail items. The brief demonstrations of their frameworks are given in (a) and (b) of Figure 1. Despite their success, two critical limitations remain unresolved: (i) their undifferentiated emphasis on tail items introduces session-irrelevant noise (e.g., ”clothing” for a session consists of books.), as not all tail items align with session-specific user requirements, resulting in the degradation of recommendation accuracy, and (ii) they lack explicit supervisory signals for long-tail objectives, thus relying on indirect optimization via cross-entropy loss. Crucially, such augmentation and re-ranking strategies often conflict with the cross-entropy optimization objective due to the inclusion of potential session-irrelevant items, resulting in a “see-saw” effect (Wang2021Denoising; Wei2024LLMRec). To address these flaws, our work revolves around two key innovations: (i) the effective discrimination of noise, restricting the consideration of long-tail issues to session-relevant tail items, and (ii) the introduction of explicit long-tail supervisory signals to concurrently improve the long-tail and accuracy performance.
@@
For the noise discrimination, given that interaction sessions are driven by user intent (Li23Intention-aware; Wang2024HearInt), we employ intent modeling to capture the overarching preference of the anonymous user. Previous work primarily derives user intents from restricted sequential segments (e.g., sliding windows) or semantically clustered items within individual sessions (Wang19Modeling; Zhang23Efficiently; Choi2024MiaSRec; Wang2024HearInt), but suffer from unreliable intent extraction due to noise interference and neglect cross-session intent consistency (Choi2024Intent; Wang2024HearInt). Therefore, we propose the hybrid intent, which captures the user preference through attribute consistency (e.g., commodity categories, music genres) and item co-occurrence patterns, as shown in (a) of Figure 2. Following this, we assign target and noise intents to each session to enable the discrimination of session-irrelevant items, as shown in (b) of Figure 2.
@@
For the long-tail supervisory signals, since the long-tail issue stems from the disparity in embedding distributions between head and tail items, resulting in discrepancies in their similarity to user embeddings (Yin2012LongTail; Gupta19NISER), we propose explicit constrains on these similarities during the training process to provide direct supervised signal. Specifically, to address the distribution inconsistency between head and tail items, we align their similarity scores to each session through a novel constraint objective. This constraint is termed the Constraint for Long-tail, which operates exclusively on session-relevant items (i.e., items belong to the target intents). Furthermore, to ensure the recommendation accuracy, we introduce an additional Constraint for Accuracy that explicitly enlarges the similarity discrepancy between sessions and session-irrelevant items (i.e., items belong to noise intents) during the training process. The mutual independence of target and noise intents ensures that the two constraints are not conflicting. The brief framework of constraints is given in Figure 1 (c).
@@
Augment-based Approaches. This technical route primarily focuses on enhancing the embeddings of tail items or emphsizing the significance of tail items when generating the session embeddings. LOAM (Yang2023LOAM) enhances tail items and sessions through the Niche-Walk Augmentation and Tail Session Mixup. GALORE (Luo2023Improving) introduces a graph augmentation approach to enhance the edge of tail items in the interaction graph. GUME (Lin2024GUME) employs the graphs and user modalities enhancement. MelT (Kim2023MELT) employs mutual enhancement of tail users and items, which jointly mitigates the long-tail issue. Additionally, some approaches have explored the useage of large language models (LLMs). LLM-ESR (Liu24LLMESR) utilizes the semantic embeddings derived from LLMs to enhance the tail items.
Rerank-based Approaches. This technical route aims to infer the distribution of tail and head items from interaction sessions, thereby enabling direct adjustment of recommendation results. TailNet (Liu2020TailNet) introduces a preference mechanism to predict the adjustment index of head and tail items. CSBR (Chen2023CSBR) proposes two additional training objectives: distribution prediction and distribution alignment to calibrate the recommendation results. LAP-SR (Peng2024LAPSR) adjusts the weight scores of recommended items based on the long-tail items and the intra-session similarity.
@@
Let $V=\{v_{1},v_{2},\dots,v_{m}\}$ represent the set of all unique items, where $m$ is their total counts. An anonymous session is represented as $S^{u}=\{v^{u}_{1},v^{u}_{2},\dots,v^{u}_{l}\}$, where $u$ is the session ID, $l$ is the length of the interaction session, and $v^{u}_{t}\in V$ (0 ¡ t ¡ $l$) is the item ID which is interacted at timestep $t$. In this paper, all symbols in bold represent the vector embeddings. For example, in $S^{u}=\{\textbf{v}^{u}_{1},\textbf{v}^{u}_{2},\dots,\textbf{v}^{u}_{l}\}$, $\textbf{v}^{u}_{t}\in\mathbb{R}^{d}$ represents the vector embedding of item $v^{u}_{t}$. Given a session $S^{u}$, the task in session-based recommendation is to predict the next-interacted item $v^{u}_{l+1}$ (i.e., the ground truth item). According to the Pareto principle (Bos1986Analysis), the top 20% of items with the highest frequency of occurrence are considered to be head items, while the remainings are tail items.
@@
Following the extraction of hybrid intent embeddings $\textbf{C}=\{\textbf{c}_{1},\textbf{c}_{2},...,\textbf{c}_{n}\}$, the subsequent objective involves imposing constraints on the learning process of session embeddings. Given the session embedding $\textbf{S}^{u}$ learned by traditional SBR models such as STAMP (liu18Stamp) or SRGNN (Wu19Srgnn), our next aim is to construct supervisory signals regarding the long-tail performance and recommendation accuracy. The demonstration of these supervisory signals are given in Figure 3.
@@
The long-tail problem emerges due to the pronounced disparity in session-item similarity between tail and head items, as documented in previous research (Yin2012LongTail; Liu20Longtail). Based on this observation, we propose a novel constraint: minimizing the variance of similarity scores between sessions and items belonging to the target intent. This constraint can reduce the divergence in similarity distributions between session-to-head and session-to-tail, thereby promoting more balanced recommendation performance. Formally, the constraint is defined as follows:
@@
$\begin{split}\min~~&\mathcal{L}_{c}=\sum_{S^{u}\in\mathcal{B}}\text{log}\frac{\text{exp}(d(\textbf{S}^{u},\textbf{c}^{u}))}{\underset{c^{v}\in\mathcal{\hat{C}}^{u}}{\sum}\text{exp}(d(\textbf{S}^{u},\textbf{c}^{v}))},~~\\ \text{s.t.}~~&\text{Var}_{c^{v}\in\mathcal{\hat{C}}^{u}}(d(\textbf{S}^{u},\textbf{c}^{v}))<\eta,\end{split}$ (9)
where exp(x) is leveraged to amplify the difference between the target and noise intents. To further minimize the effect of the noise intents, we give another theorem:
@@
$\min$~~ $\mathcal{L}_{c}=\sum_{S^{u}\in\mathcal{B}}\text{log}\frac{\text{exp}(d(\textbf{S}^{u},\textbf{c}^{u})/\sigma)}{\underset{c^{v}\in\mathcal{\hat{C}}^{u}}{\sum}\text{exp}(d(\textbf{S}^{u},\textbf{c}^{v})/\sigma)},$ (11)
@@
In addition, previous research has found that cosine similarity can achieve better alignment and uniformity of embeddings (Wang2020Understanding). Therefore, we adopt cosine similarity instead of Euclidean distance. The final training objective of the intent constraint loss (ICLoss) is formulated as:
$\begin{split}\min~~\mathcal{L}_{c}&=-\sum_{S^{u}\in\mathcal{B}}\text{log}\frac{\text{exp}(\text{cos}(\textbf{S}^{u},\textbf{c}^{u})/\sigma)}{(1+\lambda p^{u})\underset{c^{v}\in\mathcal{\hat{C}}^{u}}{\sum}\text{exp}(\text{cos}(\textbf{S}^{u},\textbf{c}^{v})/\sigma)},\end{split}$ (14)
where $\lambda$ is the hyper-parameter that controls penalty , and $p^{u}$ is rescaled within (0,1). The equivalence of Euclidean distance and cosine similarity is ensured by the $L_{2}$ normalization of Equation (5).
@@
Base Models and Competitors. To demonstrate the effectiveness of our proposed HID, we select some well-known SBR models from both sequential approaches (GRU4Rec (Hidasi16gru4rec), STAMP (liu18Stamp)) and graphic approaches (SR-GNN (Wu19Srgnn), GCE-GNN (Wang20Gcegnn)) as the base SBR models. Apart from the above base SBR models, we also introduce TailNet (Liu2020TailNet), CSBR (Chen2023CSBR), LOAM (Yang2023LOAM), LAP-SR (Peng2024LAPSR) as the plug-and-play long-tail competitors.
Metrics. To evaluate the recommendation accuracy and long-tail performance, we employ three widely used accuracy metrics, including the HR@K, and MRR@K. Following previous works on long-tail issue (Himan19Managing; Liu2020TailNet; Yang2023LOAM), we introduce some well-known long-tail metrics, including tHR@K, tMRR@K, tCov@K, and Tail@K.
@@
To investigate our proposed method, we construct two variants of our proposed method which are the HID w/o HI (i.e., Hybrid Intent) where the hybrid intent is substituted with the commonly used intent definition based on the last few items (3 in this experiment, and average pooling is adopted to aggregate them) of each session (Zhang23Efficiently), and the HID w/o FC (i.e., Flexible Coefficient) where the flexible coefficient $\sigma$ is dropped. Experiments are demonstrated in Table 2. Overall, both HID w/o HI and HID w/o FC exhibit performance degradation compared to HID across the two SBR models and datasets. Removing HI impacts diversity more, while removing FC affects accuracy more, consistent with our prior analysis. Furthermore, the hybrid intents have greater impact on Tmall than on Diginetica/Retailrocket, as its longer sessions exhibit more frequent intent shifts, making target intent modeling crucial.
@@ D   Experimental Details
Following (Xia21Cotrec), we conduct preprocessing steps over each dataset. Specifically, sessions with a length of 1 and items that appeared fewer than 5 times are excluded. Similar to (Wang20Gcegnn), we set the sessions of last week (i.e., latest data) as the test data, and the remaining historical data for training. Additionally, we use a session splitting preprocess method to augment session $S=\{s_{1},s_{2},...,s_{n}\}$ in these datasets, and generate multiple sessions with corresponding labels $([s_{1},s_{2}];s_{3}),([s_{1},s_{2},s_{3}];s_{4}),...,([s_{1},s_{2},...,s_{n-1}];s_{n})$. The statistics of the datasets are presented in Table 1.

### v2

@@ Introduction
Figure 1: Comparison between the our proposed HID and previous work. (a) illustrates the design of HID, where t. and n. denotes target and noise items for session $S^{u}$, respectively; (c) and (d) demonstrate the frameworks of previous long-tail approaches. (b) evaluates the accuracy (i.e., HR@20) and long-tail performance (i.e., tCov@20) of the base SBR model GRU4Rec (Hidasi et al. 2016) and GRU4Rec + long-tail approaches on Tmall dataset.
@@
Session-based recommendation (SBR) addresses information overload by predicting the next item from short-term interactions, particularly in privacy-sensitive scenarios lacking long-term user profiles (Li et al. 2025; Latifi et al. 2021). While deep learning methods in SBR (e.g., deep sequential models (Hidasi et al. 2016; Li et al. 2017; Liu et al. 2018; Yuan et al. 2021; Hou et al. 2022) and deep graphic models (Wu et al. 2019; Qiu et al. 2019; Wang et al. 2020; Xia et al. 2021b, a; Pan et al. 2020)) can effectively model item correlations, their model-centric focus overlooks inherent data biases. A key challenge is the long-tail distribution in recommendation data (Sundaresan 2011; Yang et al. 2023; Liu and Zheng 2020a), where a small number of high-exposure items (i.e., head items) dominate the model’s attention, while a significantly larger number of low-exposure items (i.e., tail items) are often disregarded. This unfair phenomenon leads to the overlooking of potentially essential but low-exposure tail items, limiting the diversity of recommendations (Turgut et al. 2023; Yin et al. 2024; Lee et al. 2024). Besides, the long-tail distribution causes the model to be more inclined to recommend head items, resulting in a vicious cycle.
@@
Previous advancements in long-tail SBR focus on developing plugins that seamlessly integrate with existing SBR models, emphasizing the significance of tail items to mitigate the long-tail issue (Liu and Zheng 2020a; Chen et al. 2023; Yang et al. 2023; Peng and Zhou 2024). Broadly, long-tail SBR approaches fall into two categories: (i) Augment-based approaches, which employ augmentation strategies to refine the tail item embeddings or session embeddings (Yang et al. 2023; Kim et al. 2023; Huang et al. 2024; Liu et al. 2024), and (ii) Rerank-based approaches, which predict head/tail item distributions based on interaction sessions and directly modify the final ranking results (Liu and Zheng 2020a; Chen et al. 2023; Peng and Zhou 2024). Both approaches consistently emphasize the significance of tail items. The brief demonstrations of their frameworks are given in (a) and (b) of Figure 1. Despite their success, two critical limitations remain unresolved: (i) their undifferentiated emphasis on tail items introduces session-irrelevant noise (e.g., ”clothing” for a session consists of books.), as not all tail items align with session-specific user requirements, resulting in the degradation of recommendation accuracy, and (ii) they lack explicit supervisory signals for long-tail objectives, thus relying on indirect optimization via cross-entropy loss. Crucially, such augmentation and re-ranking strategies often conflict with the cross-entropy optimization objective due to the inclusion of potential session-irrelevant items, resulting in a “see-saw” effect (Wang et al. 2021; Wei et al. 2024). To address these flaws, our work revolves around two key innovations: (i) the effective discrimination of noise, restricting the consideration of long-tail issues to session-relevant tail items, and (ii) the introduction of explicit long-tail supervisory signals to concurrently improve the long-tail and accuracy performance.
@@
For the noise discrimination, given that interaction sessions are driven by user intent (Li et al. 2023; Wang et al. 2024), we employ intent modeling to capture the overarching preference of the anonymous user. Previous work primarily derives user intents from restricted sequential segments (e.g., sliding windows) or semantically clustered items within individual sessions (Wang et al. 2019; Zhang et al. 2023; Choi et al. 2024b; Wang et al. 2024), but suffer from unreliable intent extraction due to noise interference and neglect cross-session intent consistency (Choi et al. 2024a; Wang et al. 2024). Therefore, we propose the hybrid intent, which captures the user preference through attribute consistency (e.g., commodity categories, music genres) and item co-occurrence patterns, as shown in (a) of Figure 2. Following this, we assign target and noise intents to each session to enable the discrimination of session-irrelevant items, as shown in (b) of Figure 2.
@@
For the long-tail supervisory signals, since the long-tail issue stems from the disparity in embedding distributions between head and tail items, resulting in discrepancies in their similarity to user embeddings (Yin et al. 2012; Gupta et al. 2019), we propose explicit constrains on these similarities during the training process to provide direct supervised signal. Specifically, to address the distribution inconsistency between head and tail items, we align their similarity scores to each session through a novel constraint objective. This constraint is termed the Constraint for Long-tail, which operates exclusively on session-relevant items (i.e., items belong to the target intents). Furthermore, to ensure the recommendation accuracy, we introduce an additional Constraint for Accuracy that explicitly enlarges the similarity discrepancy between sessions and session-irrelevant items (i.e., items belong to noise intents) during the training process. The mutual independence of target and noise intents ensures that the two constraints are not conflicting. The brief framework of constraints is given in Figure 1 (c).
@@
Augment-based Approaches. This technical route primarily focuses on enhancing the embeddings of tail items or emphsizing the significance of tail items when generating the session embeddings. LOAM (Yang et al. 2023) enhances tail items and sessions through the Niche-Walk Augmentation and Tail Session Mixup. GALORE (Luo et al. 2023) introduces a graph augmentation approach to enhance the edge of tail items in the interaction graph. GUME (Lin et al. 2024) employs the graphs and user modalities enhancement. MelT (Kim et al. 2023) employs mutual enhancement of tail users and items, which jointly mitigates the long-tail issue. Additionally, some approaches have explored the useage of large language models (LLMs). LLM-ESR (Liu et al. 2024) utilizes the semantic embeddings derived from LLMs to enhance the tail items.
Rerank-based Approaches. This technical route aims to infer the distribution of tail and head items from interaction sessions, thereby enabling direct adjustment of recommendation results. TailNet (Liu and Zheng 2020a) introduces a preference mechanism to predict the adjustment index of head and tail items. CSBR (Chen et al. 2023) proposes two additional training objectives: distribution prediction and distribution alignment to calibrate the recommendation results. LAP-SR (Peng and Zhou 2024) adjusts the weight scores of recommended items based on the long-tail items and the intra-session similarity.
@@
Let $V=\{v_{1},v_{2},\dots,v_{m}\}$ represent the set of all unique items, where $m$ is their total counts. An anonymous session is represented as $S^{u}=\{v^{u}_{1},v^{u}_{2},\dots,v^{u}_{l}\}$, where $u$ is the session ID, $l$ is the length of the interaction session, and $v^{u}_{t}\in V$ (0 ¡ t ¡ $l$) is the item ID which is interacted at timestep $t$. In this paper, all symbols in bold represent the vector embeddings. For example, in $S^{u}=\{\textbf{v}^{u}_{1},\textbf{v}^{u}_{2},\dots,\textbf{v}^{u}_{l}\}$, $\textbf{v}^{u}_{t}\in\mathbb{R}^{d}$ represents the vector embedding of item $v^{u}_{t}$. Given a session $S^{u}$, the task in session-based recommendation is to predict the next-interacted item $v^{u}_{l+1}$ (i.e., the ground truth item). According to the Pareto principle (Box and Meyer 1986), the top 20% of items with the highest frequency of occurrence are considered to be head items, while the remainings are tail items.
@@
Following the extraction of hybrid intent embeddings $\textbf{C}=\{\textbf{c}_{1},\textbf{c}_{2},...,\textbf{c}_{n}\}$, the subsequent objective involves imposing constraints on the learning process of session embeddings. Given the session embedding $\textbf{S}^{u}$ learned by traditional SBR models such as STAMP (Liu et al. 2018) or SRGNN (Wu et al. 2019), our next aim is to construct supervisory signals regarding the long-tail performance and recommendation accuracy. The demonstration of these supervisory signals are given in Figure 3.
@@
The long-tail problem emerges due to the pronounced disparity in session-item similarity between tail and head items, as documented in previous research (Yin et al. 2012; Liu and Zheng 2020b). Based on this observation, we propose a novel constraint: minimizing the variance of similarity scores between sessions and items belonging to the target intent. This constraint can reduce the divergence in similarity distributions between session-to-head and session-to-tail, thereby promoting more balanced recommendation performance. Formally, the constraint is defined as follows:
@@
$\begin{split}\min~~&\mathcal{L}_{c}=\sum_{S^{u}\in\mathcal{B}}\text{log}\frac{\text{exp}(d(\textbf{S}^{u},\textbf{c}^{u}))}{\text{exp}(d(\textbf{S}^{u},\textbf{c}^{u}))+\underset{c^{v}\in\mathcal{\hat{C}}^{u}}{\sum}\text{exp}(d(\textbf{S}^{u},\textbf{c}^{v}))},~~\\ \text{s.t.}~~&\text{Var}_{c^{v}\in\mathcal{\hat{C}}^{u}}(d(\textbf{S}^{u},\textbf{c}^{v}))<\eta,\end{split}$ (9)
where exp(x) is leveraged to amplify the difference between the target and noise intents, $\text{exp}(d(\textbf{S}^{u},\textbf{c}^{u}))$ is incorporated into the denominator to stabilize the loss range and mitigate the impact of the number of negative samples on the loss scale. To further minimize the effect of the noise intents, we give another theorem:
@@
$\min~~\mathcal{L}_{c}=$ (11)
$\sum_{S^{u}\in\mathcal{B}}\text{log}\frac{\text{exp}(d(\textbf{S}^{u},\textbf{c}^{u})/\sigma)}{\text{exp}(d(\textbf{S}^{u},\textbf{c}^{u})/\sigma)+\underset{c^{v}\in\mathcal{\hat{C}}^{u}}{\sum}\text{exp}(d(\textbf{S}^{u},\textbf{c}^{v})/\sigma)},$ (12)
@@
In addition, previous research has found that cosine similarity can achieve better alignment and uniformity of embeddings (Wang and Isola 2020). Therefore, we adopt cosine similarity instead of Euclidean distance. The final training objective of the intent constraint loss (ICLoss) is formulated as:
$\begin{split}\min~~\mathcal{L}_{c}=-\sum_{S^{u}\in\mathcal{B}}\text{log}\frac{\mathbf{X}}{(1+\lambda p^{u})(\mathbf{X}+\mathbf{Y})},\end{split}$ (15)
where $\lambda$ is the hyper-parameter that controls penalty, $\mathbf{X}$ is $\text{exp}(\text{cos}(\textbf{S}^{u},\textbf{c}^{u})/\sigma)$, $\mathbf{Y}$ is $\underset{c^{v}\in\mathcal{\hat{C}}^{u}}{\sum}\text{exp}(\text{cos}(\textbf{S}^{u},\textbf{c}^{v})/\sigma)$, and $p^{u}$ is rescaled within (0,1). The equivalence of Euclidean distance and cosine similarity is ensured by the $L_{2}$ normalization of Equation (5).
@@
Base Models and Competitors. To demonstrate the effectiveness of our proposed HID, we select some well-known SBR models from both sequential approaches (GRU4Rec (Hidasi et al. 2016), STAMP (Liu et al. 2018)) and graphic approaches (SR-GNN (Wu et al. 2019), GCE-GNN (Wang et al. 2020)) as the base SBR models. Apart from the above base SBR models, we also introduce TailNet (Liu and Zheng 2020a), CSBR (Chen et al. 2023), LOAM (Yang et al. 2023), LAP-SR (Peng and Zhou 2024) as the plug-and-play long-tail competitors.
Metrics. To evaluate the recommendation accuracy and long-tail performance, we employ three widely used accuracy metrics, including the HR@K, and MRR@K. Following previous works on long-tail issue (Abdollahpouri et al. 2019; Liu and Zheng 2020a; Yang et al. 2023), we introduce some well-known long-tail metrics, including tHR@K, tMRR@K, tCov@K, and Tail@K.
@@
To investigate our proposed method, we construct two variants of our proposed method which are the HID w/o HI (i.e., Hybrid Intent) where the hybrid intent is substituted with the commonly used intent definition based on the last few items (3 in this experiment, and average pooling is adopted to aggregate them) of each session (Zhang et al. 2023), and the HID w/o FC (i.e., Flexible Coefficient) where the flexible coefficient $\sigma$ is dropped. Experiments are demonstrated in Table 2. Overall, both HID w/o HI and HID w/o FC exhibit performance degradation compared to HID across the two SBR models and datasets. Removing HI impacts diversity more, while removing FC affects accuracy more, consistent with our prior analysis. Furthermore, the hybrid intents have greater impact on Tmall than on Diginetica/Retailrocket, as its longer sessions exhibit more frequent intent shifts, making target intent modeling crucial.
@@ D   Experimental Details
Following (Xia et al. 2021a), we conduct preprocessing steps over each dataset. Specifically, sessions with a length of 1 and items that appeared fewer than 5 times are excluded. Similar to (Wang et al. 2020), we set the sessions of last week (i.e., latest data) as the test data, and the remaining historical data for training. Additionally, we use a session splitting preprocess method to augment session $S=\{s_{1},s_{2},...,s_{n}\}$ in these datasets, and generate multiple sessions with corresponding labels $([s_{1},s_{2}];s_{3}),([s_{1},s_{2},s_{3}];s_{4}),...,([s_{1},s_{2},...,s_{n-1}];s_{n})$. The statistics of the datasets are presented in Table 1.